# KNN-MMD: Cross Domain Wireless Sensing via Local Distribution Alignment

### v4

@@ IV Methodology
Figure 6: Model Structure: We use the ResNet18 [24] as the feature extractor and a three-layer MLP as the predictor.
@@ IV-B Local Distribution Alignment
However, it is often difficult to find a kernel function that can realize the upper bound in practice. Therefore, Gretton et al. [43] proposed MK-MMD, which uses multiple kernel functions to approximate the upper bound, as:
@@ IV-C Model Structure
Our model architecture is illustrated in Fig. 6, which primarily consists of three components: a batch normalization layer, a feature extractor, and a final classifier. The model takes the standard CSI sequence as input, where the signal is first standardized over the time dimension, a method shown to be effective for CSI sensing model training [4]. In the network, the batch normalization layer helps reduce the domain gap between different samples, facilitating the training process. We then select the pre-trained ResNet18 [24] as the feature extractor, as the Wi-Fi signal spectrum exhibits high similarity to images. Convolutional networks, which effectively extract local features, have demonstrated excellent performance in CSI classification [22, 2, 44]. Furthermore, inspired by transfer learning, the pre-trained parameters from image data provide a strong starting point for Wi-Fi sensing [45]. Finally, we employ a multilayer perceptron (MLP) to process the embeddings for classification. The detailed description is as follows:
@@ IV-D1 Preliminary Classification
Before applying KNN for preliminary classification, we need to flatten each sample to one-dimensional data, denoted as $x^{\prime}\in R^{ls}$. Subsequently, we construct the target domain data $X^{\prime target}=[X^{\prime support},X^{\prime test}]$. However, the data dimension $ls$ is often too high, exceeding a thousand, which may lead to a high computational burden for KNN. Therefore, prior to using KNN, we aim to reduce the data dimension in $X^{\prime target}$ via UMAP [23].
UMAP primarily involves four steps: constructing the nearest neighbor graph, approximating a fuzzy-simplicial set, optimizing the low-dimensional embedding using stochastic gradient descent, and finalizing the embedding after convergence. Since UMAP is not the primary focus of our research, we directly utilize the function implemented by the UMAP package in Python. For more details, interested readers can refer to the UMAP paper [23].
After performing dimension reduction, we utilize the KNN algorithm to classify the testing set based on the support set. For more detailed information on the KNN algorithm, please refer to the paper [25]. Briefly, for each sample in the testing set, we identify the nearest $k$ samples from the support set using Euclidean distance as the distance measure. Subsequently, we determine the most frequent label among the identified neighbors as the classification result. In the case of a tie, we select the label of the nearest neighbor as the classification result.
@@ IV-D2 Training Network with Local Alignment
11:  $l^{train}\leftarrow\mathcal{L}(\hat{Y},Y^{train})$ $\triangleright$ Eq. 11
@@ V-A Dataset Description
In our experiment, we utilize two datasets: the publicly available WiGesture dataset [4] and a novel dataset named WiFall, which is proposed along with this paper. Both datasets are collected using the ESP32-S3 device, which operates in the 2.4 GHz Wi-Fi band-the default central frequency supported by the hardware and share the same data format, with a sample rate of 100Hz, 1 antenna, and 52 subcarriers. For our experiment, we divide the data into 1-second samples, resulting in each sample having a length of 100 data points.
@@ V-B Experiment Setup
Hidden Dimension of UMAP [23] 128
Hidden Dimension of ResNet [24] 64
Neighbor Number in KNN [25] : $k$ 1
@@
To evaluate the efficiency of our method, we compare it with several representative cross-domain methods in the fields of Wi-Fi sensing and machine learning. For fairness, we have made every effort to replicate other works according to the descriptions provided in their respective papers. For any missing settings in these papers, we adopt the same settings as our method. In the field of Wi-Fi sensing, most works are based on the feature learning (learning-based) paradigm. We choose the four works mentioned before for comparison: Ding et al. [18], AutoFi [19], Yang et al. [20], and CrossFi [22], which represent different powerful feature extractors and similarity calculation techniques. Even some works combined with DAL methods, such as Yang et al. [20] and CrossFi [22], which utilize the MK-MMD loss, do not rely solely on it. To further illustrate the performance of the DAL methods, we have chosen several popular methods in machine learning, including adversarial learning-based DANN [31] and ADDA [46], as well as feature-space alignment methods such as MK-MMD [43] and GFK [29]. Since these methods do not depend on specific network structures, we use the same ResNet architecture as in this paper. Additionally, we employ the same kernel functions in MK-MMD as in our KNN-MMD method. Furthermore, we also compare the methods of Tian et al. [39] and EEG [40], which also discuss local alignment methods, with detailed descriptions provided in Section III. All these DAL methods fall under the zero-shot category, meaning they only require a labeled training set from the source domain and an unlabeled testing set, without the necessity for a labeled support set.
@@ V-C Experiment Result
First, we present the experimental results in the one-shot and zero-shot scenarios as shown in Table V. Although ResNet18 [24] is not specifically designed for cross-domain scenarios, it is commonly used as a feature extractor in many related methods. Therefore, we include the results of ResNet in the in-domain scenario and zero-shot scenario as benchmarks. It can be seen that our KNN-MMD method shows excellent performance in each task, particularly outperforming all other methods in gesture recognition and action recognition tasks.
@@
Method Scenario Gesture Recognition People Identification Fall Detection Action Recognition Average Accuracy ResNet18[24] in-domain 80.75% 86.75% 91.88% 70.50% 82.47% zero-shot 40.84% 70.50% 59.86% 26.00% 49.30% Siamese[21] one-shot 70.40% 82.87% 60.62% 38.95% 63.21% AutoFi (MLP-based)[19] one-shot 24.62% 24.71% 50.88% 23.59% 30.95% AutoFi (CNN-based)[19] one-shot 27.05% 36.14% 48.05% 26.95% 34.55% Yang et al.[15] one-shot 67.21% 74.22% 59.75% 48.52% 62.43% Ding et al.[47] one-shot 39.14% 70.94% 61.56% 30.37% 50.50% CrossFi[22] one-shot 91.72% 93.01% 80.93% 49.62% 78.82% KNN [25] one-shot 83.02% 82.67% 49.63% 46.87% 65.55% KNN-MMD (Ours) one-shot 93.26% 81.84% 77.62% 75.30% 82.01% Ablation Study one-shot 69.87% 73.78% 84.03% 74.06% 75.44% MK-MMD[43] zero-shot 40.36% 66.47% 72.26% 43.72% 55.70% DANN[31] zero-shot 41.41% 67.18% 74.06% 35.99% 54.66% ADDA[46] zero-shot 42.71% 65.43% 62.81% 36.08% 51.76% GFK+KNN[29] zero-shot 30.79% 51.50% 53.72% 34.17% 42.55% CrossFi[22] zero-shot 64.81% 72.79% 74.38% 40.46% 63.11% Tian et al.[39] zero-shot 68.13% 55.86% 61.72% 42.10% 56.95% EEG[40] zero-shot 59.75% 64.63% 69.53% 42.15% 59.02%
@@ V-D1 Metric-based Method
For metric-based methods like KNN [25], the biggest problem is that they are heavily influenced by the quality of the support set. As shown in Table VI, we reduced the data dimension in the target domain to $d$ using UMAP and tested the KNN performance in an $n$-shot scenario with $k$ nearest neighbors. Each scenario was tested 3 times, and the table shows the accuracy range. It can be seen that the accuracy fluctuates greatly, which may pose challenges for the practical application of this method. In Fig. 9, it can be seen that our KNN-MMD method does not exhibit such large fluctuations. Moreover, in Table VI, it’s noticeable that our KNN-MMD is not very sensitive to the hyper-parameters.
TABLE VI: Performance of KNN [25] and Our KNN-MMD: $n$ denotes the number of shots, $k$ denotes the number of neighbors in KNN, and $d$ denotes the data dimension after reduction using UMAP [23].
KNN [25] KNN-MMD
@@ V-D2 Learning-based Method
For learning-based methods, the biggest problem is that we do not know when to stop the training process. Fig. 10a illustrates the training process of a Siamese network [21], a traditional one-shot learning model. It can be seen that the performance change is relatively stable on the training set, but not on the testing set. However, during training, we do not have direct access to the performance on the testing set. This makes it very possible to stop the model training at an epoch that has worse performance on the testing set. Our KNN-MMD method efficiently solves this problem by using early stop according to the accuracy in the support set. In traditional learning-based methods, the support set also participates in training, so our early stop method cannot be used in them. Furthermore, as shown in Fig. 10c, the testing accuracy of our KNN-MMD also increases stably with the help of our local alignment mechanism.
(a) Siamese Network [21]
(b) MK-MMD [43]
@@ V-F Expanded Experiment
To further illustrate the efficiency of our KNN-MMD, we conduct cross-domain experiments using the WiDar3.0 dataset [48], which includes both cross-people and cross-room gesture recognition scenarios. In WiDar3.0, the Body-coordinate Velocity Profile (BVP) extracted from CSI is provided, with the dimension of $t\times 20\times 20$, where $t$ represents the time length. To standardize the data dimensions, we utilize only the first 28 frames of each sample and filter out those samples that are shorter than 28 frames. For our experiment, we select the subsets as follows: (i) Cross-People Scenario: For the cross-person gesture task, we utilize data collected on November 9, 2018, which includes six gestures: Push&Pull, sweep, clap, slide, draw, zigzag, and draw-N. We designate User 1 as the source domain and User 2 as the target domain. Both the source and target domains contain 3,000 samples. (ii) Cross-Room Scenario: For the cross-room gesture task, we utilize data collected on November 21 and November 27, 2018, from User 2, which includes six gestures: slide, draw-O, draw zigzag, draw-N, draw-triangle, and draw-rectangle. We designate Room 1 as the source domain and Room 2 as the target domain. The source domain contains 3,000 samples, while the target domain has 750 samples.
The experimental results are shown in Fig. 12, where our KNN-MMD demonstrates excellent performance in both cross-people and cross-room tasks, slightly outperforming CrossFi and significantly outperforming other benchmarks. Since AutoFi [19] only supports CSI input due to its pre-training method, we have omitted it from this section.
@@
Figure 12: Experiment Results on WiDar3.0 [48]
@@ V-F2 Influence of Model Input
In many Wi-Fi sensing studies, processed CSI signals, such as Time of Flight (ToF), Angle of Arrival (AoA), Attenuation, and Doppler Frequency Shift (DFS), are commonly used instead of the original CSI. These processed signals serve as important feature extraction methods in model-based approaches. However, this practice remains controversial in the field of deep learning. While data pre-processing methods can help extract meaningful features, they may also lead to the loss of information present in the original data. As noted in [48], DFS contains the most information about velocity distribution, offering advantages over other features. Additionally, Micro-Doppler has been successfully used in OFDM radar sensing tasks and has also shown promising results in Wi-Fi sensing [49]. To investigate this, we compare the performance of KNN-MMD on the WiGesture dataset using both original CSI signals, DFS, and Micro-Doppler. Specifically, DFS is obtained by estimating the frequency shift from the phase differences of consecutive CSI samples, which primarily captures macro motions and change. For a CSI stream denoted as $H(t)=|H(t)|e^{j\angle H(t)}$, the phase difference between consecutive samples is used to estimate the Doppler shift $f_{d}$ via: $f_{d}=\frac{1}{2\pi}\frac{\mathrm{d}(\angle H(t))}{\mathrm{d}t}$. Micro-Doppler, in contrast, is extracted by applying time-frequency analysis such as the Short-Time Fourier Transform (STFT) to the CSI series, which reveals finer-grained motion patterns. Shown as Fig. 13, the results indicate that, compared to original CSI input, DFS can improve performance in people identification but may lead to lower performance in gesture recognition. Conversely, Micro-Doppler demonstrates an opposite trend. The underlying reasons for these observations may be complex; while data pre-processing can provide prior knowledge to the network, it may also result in potential information loss from the original data. Nonetheless, the overall performance of our method does not vary significantly among different inputs, illustrating its robustness. In practice, the choice of input can be tailored to the specific task and available computational resources.
Figure 13: Experiment Results on the WiGesture Dataset [4]: The solid line, dashed line, and dotted line represent model inputs using original CSI signals, DFS, and Micro-Doppler, respectively.

### v5

@@ IV Methodology
Figure 6: Model Structure: We use the ResNet18 [ResNet] as the feature extractor and a three-layer MLP as the predictor.
@@ IV-B Local Distribution Alignment
However, it is often difficult to find a kernel function that can realize the upper bound in practice. Therefore, Gretton et al. [MK-MMD] proposed MK-MMD, which uses multiple kernel functions to approximate the upper bound, as:
@@ IV-C Model Structure
Our model architecture is illustrated in Fig. 6, which primarily consists of three components: a batch normalization layer, a feature extractor, and a final classifier. The model takes the standard CSI sequence as input, where the signal is first standardized over the time dimension, a method shown to be effective for CSI sensing model training [CSI-BERT]. In the network, the batch normalization layer helps reduce the domain gap between different samples, facilitating the training process. We then select the pre-trained ResNet18 [ResNet] as the feature extractor, as the Wi-Fi signal spectrum exhibits high similarity to images. Convolutional networks, which effectively extract local features, have demonstrated excellent performance in CSI classification [CSi-Net, fall, cai2023falldewideo]. Furthermore, inspired by transfer learning, the pre-trained parameters from image data provide a strong starting point for Wi-Fi sensing [zhu2024srcsense]. Finally, we employ a multilayer perceptron (MLP) to process the embeddings for classification. The detailed description is as follows:
@@ IV-D1 Preliminary Classification
Before applying KNN for preliminary classification, we need to flatten each sample to one-dimensional data, denoted as $x^{\prime}\in R^{ls}$. Subsequently, we construct the target domain data $X^{\prime target}=[X^{\prime support},X^{\prime test}]$. However, the data dimension $ls$ is often too high, exceeding a thousand, which may lead to a high computational burden for KNN. Therefore, prior to using KNN, we aim to reduce the data dimension in $X^{\prime target}$ via UMAP [UMAP].
UMAP primarily involves four steps: constructing the nearest neighbor graph, approximating a fuzzy-simplicial set, optimizing the low-dimensional embedding using stochastic gradient descent, and finalizing the embedding after convergence. Since UMAP is not the primary focus of our research, we directly utilize the function implemented by the UMAP package in Python. For more details, interested readers can refer to the UMAP paper [UMAP].
After performing dimension reduction, we utilize the KNN algorithm to classify the testing set based on the support set. For more detailed information on the KNN algorithm, please refer to the paper [KNN]. Briefly, for each sample in the testing set, we identify the nearest $k$ samples from the support set using Euclidean distance as the distance measure. Subsequently, we determine the most frequent label among the identified neighbors as the classification result. In the case of a tie, we select the label of the nearest neighbor as the classification result.
@@ IV-D2 Training Network with Local Alignment
11:  $l^{train}\leftarrow\mathcal{L}(\hat{Y},Y^{train};X^{train},X^{support})$ $\triangleright$ Eq. 11
@@ V-A Dataset Description
In our experiment, we utilize two datasets: the publicly available WiGesture dataset [CSI-BERT] and a novel dataset named WiFall, which is proposed along with this paper. Both datasets are collected using the ESP32-S3 device, which operates in the 2.4 GHz Wi-Fi band-the default central frequency supported by the hardware and share the same data format, with a sample rate of 100Hz, 1 antenna, and 52 subcarriers. For our experiment, we divide the data into 1-second samples, resulting in each sample having a length of 100 data points.
@@ V-B Experiment Setup
Hidden Dimension of UMAP [UMAP] 128
Hidden Dimension of ResNet [ResNet] 64
Neighbor Number in KNN [KNN] : $k$ 1
@@
To evaluate the efficiency of our method, we compare it with several representative cross-domain methods in the fields of Wi-Fi sensing and machine learning. For fairness, we have made every effort to replicate other works according to the descriptions provided in their respective papers. For any missing settings in these papers, we adopt the same settings as our method. In the field of Wi-Fi sensing, most works are based on the feature learning (learning-based) paradigm. We choose the four works mentioned before for comparison: Ding et al. [ding2021wi], AutoFi [AutoFi], Yang et al. [yang2019learning], and CrossFi [CSi-Net], which represent different powerful feature extractors and similarity calculation techniques. Even some works combined with DAL methods, such as Yang et al. [yang2019learning] and CrossFi [CSi-Net], which utilize the MK-MMD loss, do not rely solely on it. To further illustrate the performance of the DAL methods, we have chosen several popular methods in machine learning, including adversarial learning-based DANN [DANN] and ADDA [ADDA], as well as feature-space alignment methods such as MK-MMD [MK-MMD] and GFK [GFK]. Since these methods do not depend on specific network structures, we use the same ResNet architecture as in this paper. Additionally, we employ the same kernel functions in MK-MMD as in our KNN-MMD method. Furthermore, we also compare the methods of Tian et al. [KNN-MMD] and EEG [EEG], which also discuss local alignment methods, with detailed descriptions provided in Section III. All these DAL methods fall under the zero-shot category, meaning they only require a labeled training set from the source domain and an unlabeled testing set, without the necessity for a labeled support set.
@@ V-C Experiment Result
First, we present the experimental results in the one-shot and zero-shot scenarios as shown in Table V. Although ResNet18 [ResNet] is not specifically designed for cross-domain scenarios, it is commonly used as a feature extractor in many related methods. Therefore, we include the results of ResNet in the in-domain scenario and zero-shot scenario as benchmarks. It can be seen that our KNN-MMD method shows excellent performance in each task, particularly outperforming all other methods in gesture recognition and action recognition tasks.
@@
Method Scenario Gesture Recognition People Identification Fall Detection Action Recognition Average Accuracy ResNet18[ResNet] in-domain 80.75% 86.75% 91.88% 70.50% 82.47% zero-shot 40.84% 70.50% 59.86% 26.00% 49.30% Siamese[siamese] one-shot 70.40% 82.87% 60.62% 38.95% 63.21% AutoFi (MLP-based)[AutoFi] one-shot 24.62% 24.71% 50.88% 23.59% 30.95% AutoFi (CNN-based)[AutoFi] one-shot 27.05% 36.14% 48.05% 26.95% 34.55% Yang et al.[SN-MMD] one-shot 67.21% 74.22% 59.75% 48.52% 62.43% Ding et al.[CNN-based] one-shot 39.14% 70.94% 61.56% 30.37% 50.50% CrossFi[CSi-Net] one-shot 91.72% 93.01% 80.93% 49.62% 78.82% KNN [KNN] one-shot 83.02% 82.67% 49.63% 46.87% 65.55% KNN-MMD (Ours) one-shot 93.26% 81.84% 77.62% 75.30% 82.01% Ablation Study one-shot 69.87% 73.78% 84.03% 74.06% 75.44% MK-MMD[MK-MMD] zero-shot 40.36% 66.47% 72.26% 43.72% 55.70% DANN[DANN] zero-shot 41.41% 67.18% 74.06% 35.99% 54.66% ADDA[ADDA] zero-shot 42.71% 65.43% 62.81% 36.08% 51.76% GFK+KNN[GFK] zero-shot 30.79% 51.50% 53.72% 34.17% 42.55% CrossFi[CSi-Net] zero-shot 64.81% 72.79% 74.38% 40.46% 63.11% Tian et al.[KNN-MMD] zero-shot 68.13% 55.86% 61.72% 42.10% 56.95% EEG[EEG] zero-shot 59.75% 64.63% 69.53% 42.15% 59.02%
@@ V-D1 Metric-based Method
For metric-based methods like KNN [KNN], the biggest problem is that they are heavily influenced by the quality of the support set. As shown in Table VI, we reduced the data dimension in the target domain to $d$ using UMAP and tested the KNN performance in an $n$-shot scenario with $k$ nearest neighbors. Each scenario was tested 3 times, and the table shows the accuracy range. It can be seen that the accuracy fluctuates greatly, which may pose challenges for the practical application of this method. In Fig. 9, it can be seen that our KNN-MMD method does not exhibit such large fluctuations. Moreover, in Table VI, it’s noticeable that our KNN-MMD is not very sensitive to the hyper-parameters.
TABLE VI: Performance of KNN [KNN] and Our KNN-MMD: $n$ denotes the number of shots, $k$ denotes the number of neighbors in KNN, and $d$ denotes the data dimension after reduction using UMAP [UMAP].
KNN [KNN] KNN-MMD
@@ V-D2 Learning-based Method
For learning-based methods, the biggest problem is that we do not know when to stop the training process. Fig. 10a illustrates the training process of a Siamese network [siamese], a traditional one-shot learning model. It can be seen that the performance change is relatively stable on the training set, but not on the testing set. However, during training, we do not have direct access to the performance on the testing set. This makes it very possible to stop the model training at an epoch that has worse performance on the testing set. Our KNN-MMD method efficiently solves this problem by using early stop according to the accuracy in the support set. In traditional learning-based methods, the support set also participates in training, so our early stop method cannot be used in them. Furthermore, as shown in Fig. 10c, the testing accuracy of our KNN-MMD also increases stably with the help of our local alignment mechanism.
(a) Siamese Network [siamese]
(b) MK-MMD [MK-MMD]
@@ V-F Expanded Experiment
To further illustrate the efficiency of our KNN-MMD, we conduct cross-domain experiments using the WiDar3.0 dataset [zhang2021widar3], which includes both cross-people and cross-room gesture recognition scenarios. In WiDar3.0, the Body-coordinate Velocity Profile (BVP) extracted from CSI is provided, with the dimension of $t\times 20\times 20$, where $t$ represents the time length. To standardize the data dimensions, we utilize only the first 28 frames of each sample and filter out those samples that are shorter than 28 frames. For our experiment, we select the subsets as follows: (i) Cross-People Scenario: For the cross-person gesture task, we utilize data collected on November 9, 2018, which includes six gestures: Push&Pull, sweep, clap, slide, draw, zigzag, and draw-N. We designate User 1 as the source domain and User 2 as the target domain. Both the source and target domains contain 3,000 samples. (ii) Cross-Room Scenario: For the cross-room gesture task, we utilize data collected on November 21 and November 27, 2018, from User 2, which includes six gestures: slide, draw-O, draw zigzag, draw-N, draw-triangle, and draw-rectangle. We designate Room 1 as the source domain and Room 2 as the target domain. The source domain contains 3,000 samples, while the target domain has 750 samples.
The experimental results are shown in Fig. 12, where our KNN-MMD demonstrates excellent performance in both cross-people and cross-room tasks, slightly outperforming CrossFi and significantly outperforming other benchmarks. Since AutoFi [AutoFi] only supports CSI input due to its pre-training method, we have omitted it from this section.
@@
Figure 12: Experiment Results on WiDar3.0 [zhang2021widar3]
@@ V-F2 Influence of Model Input
In many Wi-Fi sensing studies, processed CSI signals, such as Time of Flight (ToF), Angle of Arrival (AoA), Attenuation, and Doppler Frequency Shift (DFS), are commonly used instead of the original CSI. These processed signals serve as important feature extraction methods in model-based approaches. However, this practice remains controversial in the field of deep learning. While data pre-processing methods can help extract meaningful features, they may also lead to the loss of information present in the original data. As noted in [zhang2021widar3], DFS contains the most information about velocity distribution, offering advantages over other features. Additionally, Micro-Doppler has been successfully used in OFDM radar sensing tasks and has also shown promising results in Wi-Fi sensing [tang2023mdpose].
To investigate this, we compare the performance of KNN-MMD on the WiGesture dataset using both original CSI signals, DFS, and Micro-Doppler. Specifically, DFS is obtained by estimating the frequency shift from the phase differences of consecutive CSI samples, which primarily captures macro motions and change. For a CSI stream denoted as $H(t)=|H(t)|e^{j\angle H(t)}$, the phase difference between consecutive samples is used to estimate the Doppler shift $f_{d}$ via: $f_{d}=\frac{1}{2\pi}\frac{\mathrm{d}(\angle H(t))}{\mathrm{d}t}$. Considering the CSI sequence $x\in\mathbb{R}^{l\times s}$, where $l$ is the number of frames and $s$ is the number of subcarriers, we obtain the DFS feature with dimensions $\mathbb{R}^{(l-1)\times s}$. We then expand this feature to include an additional dimension for input into ResNet, resulting in dimensions of $\mathbb{R}^{1\times(l-1)\times s}$.
Micro-Doppler, in contrast, is extracted by applying time-frequency analysis such as the Short-Time Fourier Transform (STFT) to the CSI series, which reveals finer-grained motion patterns. Specifically, we set the frequency bins to $f$ to obtain a frequency feature with dimensions $\mathbb{R}^{s\times f\times t}$, where $t$ denotes the time frames after the STFT. We then treat the subcarrier dimension $s$ as the channel and input it into the ResNet for feature extraction. In the experiment, we set the window size to 60, the overlap to 58, and the number of frequency bins to 128.
Shown as Fig. 13, the results indicate that, compared to original CSI input, DFS can improve performance in people identification but may lead to lower performance in gesture recognition. Conversely, Micro-Doppler demonstrates an opposite trend. This is because DFS effectively captures the low-frequency, consistent stylistic patterns of an individual’s movement, which is stable for identity. In contrast, Micro-Doppler excels at extracting the high-frequency, detailed kinematics of a specific motion, making it superior for classifying gesture types but more sensitive to individual variations that hinder identification.
Figure 13: Experiment Results on the WiGesture Dataset: The blue lines represent the gesture recognition tasks, while the red lines indicate the people identification tasks.